A pharmacist wishes to mix a solution that is 4% minoxidil. She has on hand 70mL of a 1% solution, and she wishes to add some 6% solution to it to obtain the desired 4% solution. How much 6% solution should she add?
Answer by stanbon(75887)

A pharmacist wishes to mix a solution that is 4% minoxidil. She has on hand 70mL of a 1% solution, and she wishes to add some 6% solution to it to obtain the desired 4% solution. How much 6% solution should she add?
-----------------
Equation:
mini + mini = mini
0.01*70 + 0.06x = 0.04(70+x)
----
1*70 + 6x = 4*70 + 4x
-----
2x = 3*70
---
x = 3*35
----
x = 105 mL (amt. of 6% solution needed)
==============================
Cheers,
Stan H.
===================